OpenStudy (anonymous):

Write the standard equation of the circle center (-6,-8), that passes through (0,0)

OpenStudy (anonymous):

standard equation: \[(x - h)^2 + (y - k)^2 = r^2\] where (h, k) is the circle's center and r is the circle's radius. the radius must be the distance between (-6. -8) and (0, 0), which can be found using the distance formula. \[r = d = \sqrt{ (x_2 - x_1)^2 + (y_2 - y_1)^2 }\]